A rectangle box with a base 6 by 8 has a diagonal of length 20. Find the exact height of the box.
Answer by edjones(8007)

6^2+8^2
=36+64
=100
sqrt(100)=10 diagonal of the base of the box.
.
hypotenuse=20 given. Goes from the base of one corner to the top of the other corner.
a^2+10^2=20^2
a^2+100=400
a^2=300
a=sqrt(300)
=sqrt(100*3)
=10sqrt(3) exact height of box.
.
Ed